OpenStudy (anonymous):

Using the graph below, calculate the average rate of change for f(x) from x = 2 to x = 6. exponential function going through points 2, 2 and 6, 6 x = −4 x = −1 x = 1 x = 4 @wio @ganeshie8 @TheSmartOne Plz help me

OpenStudy (anonymous):

I think it's D

ganeshie8 (ganeshie8):

average rate of change between two points is same as the slope of line segment joining those two points

OpenStudy (anonymous):

I'm right? it has to be either A or D

ganeshie8 (ganeshie8):

(2, 2) and (6, 6) find the slope of line passing through these two points

OpenStudy (anonymous):

oh no it's 1 right?

ganeshie8 (ganeshie8):

yes average rate of change is 1

ganeshie8 (ganeshie8):

but your options make no sense

OpenStudy (anonymous):

yeah ik..

OpenStudy (anonymous):

What exponential function is the best fit for the data in the table? x f(x) 1 -4 3 -1 4 3 f(x) = 4(3)x - 1 + 4 f(x) = 4(3)x - 1 - 4 f(x) = one fourth(3)x - 1 + 4 f(x) = one fourth(3)x - 1 - 4 Can u plz help me with this one also? @ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

For an exponential function...

OpenStudy (anonymous):

i think the answer is D but nit sure..

OpenStudy (anonymous):

is it D? @wio

OpenStudy (anonymous):

Let me think about it for a moment

OpenStudy (anonymous):

\[ y=ab^x+c \]

OpenStudy (anonymous):

\[ -4=ab+c\\ -1=ab^3+c\\ 3=ab^4+c \]

OpenStudy (anonymous):

ok its D can u check this for me quickly plz i need to go to sleep in a few

OpenStudy (anonymous):

What statement correctly describes the key features of the graph of f(x) = 4(one half)x + 1 − 3? Y-intercept of (0, −1), starts down on the left, gets closer to y = −3 on the right Y-intercept of (0, −1), starts up on the left, gets closer to y = −3 on the right <-- MY ANSWER Y-intercept of (0, 1), starts up on the left, gets closer to y = −3 on the right Y-intercept of (0, 1), starts down on the left, gets closer to y = −3 on the right

OpenStudy (anonymous):

nvm ur kinda slow a little no offense im done already